Evaluate -3x^3 - 4x for x = -1.

a.) 1
b.) 7
c.) -1

Answer by ikleyn(52824)

.
My dear, substitute x = -1 instead of x into the expression -3x^3 - 4x. You will get

-3*(-1)^3 - 4*(-1) = -3*(-1) + 4 = 3 + 4 = 7.


That's all.